# A Framework of SO(3)-equivariant Non-linear Representation Learning and its Application to Electronic-Structure Hamiltonian Prediction

### v4

@@ 6.2 Results and Analysis
We list the results of DeepH-E3 and DeepH-E3+TraceGrad in Tables 1 for databases from the DeepH benchmark series, reporting the values of $MAE^{H}_{all}$, $MAE^{H}_{cha\_s}$, and $MAE^{H}_{cha\_b}$. The results of DeepH-E3 to be compared are copied from Yin et al. (2024). The results of DeepH-E3+TraceGrad are the average from 10 independent repeated experiments. Regarding the metric of $MAE^{H}_{block}$ for every Hamiltonian block, due to its large data volume, we just present its values for all databases from the DeepH series in Appendix 4. We use the same fixed random seed as adopted by DeepH-E3 for all random processes in experiments on these six databases. As a result, the standard deviation of the Hamiltonian prediction MAE across repeated experiments does not exceed $0.007$ meV for each of the six databases and is negligible.
@@
From results presented in Tables 1, we could find that the proposed TraceGrad method dramatically enhances the accuracy performance of the baseline method DeepH-E3, both on average and for challenging samples and blocks, both on the non-twisted samples and the twisted samples. Specifically, on the corresponding datasets, TraceGrad lowers down the $MAE^{H}_{all}$ and $MAE^{H}_{cha}$ of DeepH-E3 with relative ratios of up to $34\%$ and $35\%$, respectively. Furthermore, from the results included in Appendix 4, TraceGrad significantly improves the performance for the vast majority of basic blocks. Particularly, for the blocks where DeepH-E3 perform the worst, TraceGrad reduces the $MAE$ ($MAE^{H}_{cha\_b}$) by a maximum of $48\%$. The leading performance on the $MG$ and $MM$ databases prepared at three-hundred Kelvin temperature demonstrates the robustness of our method against thermal motion. The high accuracy on the $BB$, $BT$, and $BS$ databases, which have strong SOC effects, indicates our method’s strong capability to model such effects. The excellent performance on the $BG^{t}$, $BB^{t}$, $BT^{t}$, and $BS^{t}$ subsets showcases the method’s superior generalization to twisted structures, which are not present in the training data. The outstanding performance on such samples highlights the good potential for studying twist-related phenomena, a hot research topic that may bring new electrical and transport properties (Cao et al., 2018; Wang et al., 2024a; He et al., 2024). Additionally, the $BG^{t}$, $BB^{t}$, $BT^{t}$, and $BS^{t}$ subsets contain significantly larger unit cells compared to the training set (see Appendix C for statistics of their sizes), yet our method still excels on these subsets as measured by the multiple MAE metrics, demonstrating its good scalability on the sizes of atomic systems it handles.
@@
In Table 2, we present the results of QHNet and QHNet+TraceGrad under the metrics of $MAE^{H}_{all}$, $MAE^{H}_{diag}$, $MAE^{H}_{non\_diag}$, $MAE^{\epsilon}$, and $Sim(\psi)$ for the $QS$ and $QD$ databases. The results of QHNet to be compared are taken from their original paper (Yu et al., 2023a), and for the unification of MAE units, we convert the units of MAE from $10^{-6}$ Hartree ($E_{h}$) in the original paper to meV111 $27211.41E_{h}=$ 27211.4 meV. The results of QHNet+TraceGrad are the average from 10 independent repeated experiments. To ensure reproducibility, we use the same fixed random seeds as employed in QHNet for all random processes in the experiments on the $QS$ and $QD$ databases. As a result, the standard deviation of the Hamiltonian prediction MAE across repeated experiments is no greater than $0.009$ meV for both $QS$ and $QD$ and is also negligible.
@@
The results presented in Table 2 demonstrate that the proposed TraceGrad method significantly enhances the accuracy of the baseline QHNet model across all metrics on the $QS$ and $QD$ databases. Specifically, TraceGrad reduces $MAE^{H}_{all}$, $MAE^{H}_{non\_diag}$, $MAE^{H}_{diag}$, $MAE^{\epsilon}$, and $Sim(\psi)$ of QHNet with relative reductions of up to 40%, 39%, 40%, 76%, and 17%, respectively, on the corresponding databases. The significant accuracy improvements on the $QS$ database, partitioned using the ’ood’ split strategy (Yu et al., 2023a) without scale overlapping among the training, validation, and testing sets, once again demonstrate the method’s strong generalization capabilities across different scales of atomic systems. Meanwhile, performance on the $QD$ database under the ’mol’ strategy (Yu et al., 2023a), which partitions the training, validation, and testing sets with samples from completely different thermal motion trajectories, highlight our method’s robustness in generalizing to new thermal motion sequences. Furthermore, the substantial improvement in the prediction accuracy of $\epsilon$, i.e., occupied orbital energies crucial for determining electronic properties such as optical characteristics and conductivity in atomic systems, and $\psi$, i.e., the electronic wavefunctions essential for understanding electron distribution and interactions, underscores the potential values of our method for applications like material design, molecular pharmacology, and quantum computing.
We also compare the acceleration ratios brought by the QHNet model and the QHNet+TraceGrad model for the convergence of classical DFT algorithms implemented by PySCF (Sun et al., 2018), with the experimental results reported in Table 3, using the achieved ratio and error-level ratio metrics introduced in Section 6.1. As observed from this table, TraceGrad also brings significant improvements to the baseline model QHNet, notably reducing the achieved ratio and enhancing the error-level ratio, indicating its substantial potential for accelerating traditional DFT algorithms.
Due to the page limits, we present three groups of supplementary experimental results in the Appendices. The first group, included in Appendix H, is the results of ablation study. These fine-grained experiments assess the individual and combined contributions of the two core mechanisms of the proposed TraceGrad method on the six databases from the DeepH series. These mechanisms are the SO(3)-invariant trace supervision mechanism (Trace) at the label level and the gradient-based induction mechanism (Grad) at the representation layer. Results indicate that each mechanism can contribute individually to the performance. Moreover, their combination provides even better performance. The second group of experiments, as included in Appendix I, test the extra time cost caused by TraceGrad compare to the baseline model. As we adopt a light-weight implementation of the non-linear modules introduced in the TraceGrad method, experimental results show that the extra time burden is slight relative to the baseline models DeepH-E3 and QHNet, indicating the good efficiency of our approach. The third group of experiments corresponds to the synergy of our method with an approximately SO(3)-equivariant methodology (Yin et al., 2024). As approximate equivariance is not the main focus of this paper, we just include these experimental results in Appendix J. The experimental results in Appendix J still demonstrate a significant accuracy improvement brought by our approach, further supporting the generality of our methodology.
@@ 7 Conclusion
We propose a theoretical and methodological framework to tackle the issue of reconciling non-linear expressiveness with SO(3)-equivariance in deep learning frameworks for physical system modeling, through deeply investigating the mathematical connections between SO(3)-invariant and SO(3)-equivariant quantities, as well as their representations. We first constructs SO(3)-invariant quantities from SO(3)-equivariant regression targets, using them to train informative SO(3)-invariant non-linear representations. From these, SO(3)-equivariant features are derived with gradient operations, achieving non-linear expressiveness while maintaining strict SO(3)-equivariance. We apply our theory and method to the challenging electronic-structure Hamiltonian prediction tasks, achieving dramatic promotions in prediction accuracy across eight benchmark databases. Experimental results demonstrate that this approach not only improves the accuracy of Hamiltonian prediction but also significantly enhances the prediction for downstream physical quantities, and also markedly improves the acceleration ratios for traditional DFT algorithms.
Ethics Statement
This work develops a representation learning method that exhibits strong non-linear expressive capabilities while strictly adhering to SO(3) equivariance. This method has demonstrated superior accuracy in predicting electronic-structure Hamiltonians and related physical quantities, showcasing its potential to accelerate research in materials science and molecular pharmacology. While we recognize that our research area has not yet revealed direct negative social or ethical implications, several issues warrant our vigilance. Currently, although our method yields accurate predictions, the decision-making processes of deep learning systems often lack transparency, hindering a comprehensive understanding of the learning outcomes and limiting our ability to gain deeper insights. We believe it is important to investigate the interpretability of such models, particularly in terms of how they apply physical knowledge in a comprehensible way. Additionally, it is crucial to continually improve the correctness and fairness of deep learning models on this area. Ensuring high-quality and diverse training data, implementing sound model designs, and performing ongoing validation and refinement are necessary to guarantee model accuracy and the broad applicability of their results.
@@
A Background
Definition 1.
Group. A set $G$, denoted as $G=\{\dots,g,\dots\}$, equipped with a binary operation denoted as $\cdot$, is called a group if it satisfies the following four conditions:
@@ A Background
    Closure: For any two elements $f,g\in G$, $f\cdot g\in G$.
@@
    Associativity: For all elements $f,g,h\in G$, the equality $(f\cdot g)\cdot h=f\cdot(g\cdot h)$ holds.
@@
    Existence of Identity Element: There exists a unique identity element $e\in G$ such that for all $f\in G$, $e\cdot f=f\cdot e=f$.
@@
    Existence of Inverse Element: For each element $f\in G$, there exists a unique inverse element $f^{-1}\in G$ such that $f\cdot f^{-1}=f^{-1}\cdot f=e$.
Definition 2.
SO(3) Group. The special orthogonal group $\text{SO}(3)$ is the group of all $3\times 3$ orthogonal matrices with determinant 1. Formally, it can be defined as:
$\text{SO}(3)=\{\mathbf{R}\in\mathbb{R}^{3\times 3}\mid\mathbf{R}^{T}\mathbf{R}=I,\,\det(\mathbf{R})=1\}$
where $\mathbf{R}^{T}$ denotes the transpose of $\mathbf{R}$ and $I$ is the $3\times 3$ identity matrix. The elements of $\text{SO}(3)$ represent rotations in three-dimensional Euclidean space.
Definition 3.
Group Representation. A representation of a group $G$ on a tensor space $T(V)$ is a homomorphism $\rho$ from $G$ to the general linear group $GL(T(V))$, the group of all invertible linear transformations on $T(V)$. Here, $T(V)$ represents the tensor space associated with the vector space $V$, encompassing all tensors that can be formed from elements of $V$. The homomorphism $\rho$ can be formalized as:
@@
Definition 4.
SO(3) Group Representation. A representation of the special orthogonal group $\text{SO}(3)$ on a tensor space $T(V)$ is a homomorphism $\rho:\text{SO}(3)\to GL(T(V))$, where $GL(T(V))$ denotes the group of all invertible linear transformations on the tensor space $T(V)$. This representation can be expressed using Wigner-D matrices $\mathbf{D}^{l}(\mathbf{R})$, which are defined for a given degree $l$ corresponding to the representation of the rotation matrix $\mathbf{R}\in\text{SO}(3)$:
$\rho(\mathbf{R})=\mathbf{D}^{l}(\mathbf{R}).$
Definition 5.

### v5

@@ 6.2 Results and Analysis
We list the results of DeepH-E3 and DeepH-E3+TraceGrad in Table 1 for databases from the DeepH benchmark series, reporting the values of $MAE^{H}_{all}$, $MAE^{H}_{cha\_s}$, and $MAE^{H}_{cha\_b}$. The results of DeepH-E3 to be compared are copied from Yin et al. (2024). The results of DeepH-E3+TraceGrad are the average from 10 independent repeated experiments. Regarding the metric of $MAE^{H}_{block}$ for every Hamiltonian block, due to its large data volume, we just present its values for all databases from the DeepH series in Appendix G. We use the same fixed random seed as adopted by DeepH-E3 for all random processes in experiments on these six databases. As a result, for each Hamiltonian MAE result presented in Table 1, the standard deviation across independent repeated experiments does not exceed $0.007$ meV and is negligible.
@@
From results presented in Tables 1, we could find that the proposed TraceGrad method dramatically enhances the accuracy performance of the baseline method DeepH-E3, both on average and for challenging samples and blocks, both on the non-twisted samples and the twisted samples. Specifically, on the corresponding datasets, TraceGrad lowers down the $MAE^{H}_{all}$ and $MAE^{H}_{cha}$ of DeepH-E3 with relative ratios of up to $34\%$ and $35\%$, respectively. Furthermore, from the results included in Appendix G, TraceGrad significantly improves the performance for the vast majority of basic blocks. Particularly, for the blocks where DeepH-E3 perform the worst, TraceGrad reduces the $MAE$ ($MAE^{H}_{cha\_b}$) by a maximum of $48\%$. The leading performance on the $MG$ and $MM$ databases prepared at three-hundred Kelvin temperature demonstrates the robustness of our method against thermal motion. The high accuracy on the $BB$, $BT$, and $BS$ databases, which have strong SOC effects, indicates our method’s strong capability to model such effects. The excellent performance on the $BG^{t}$, $BB^{t}$, $BT^{t}$, and $BS^{t}$ subsets showcases the method’s superior generalization to twisted structures, which are not present in the training data. The outstanding performance on such samples highlights the good potential for studying twist-related phenomena, a hot research topic that may bring new electrical and transport properties (Cao et al., 2018; Wang et al., 2024a; He et al., 2024). Additionally, the $BG^{t}$, $BB^{t}$, $BT^{t}$, and $BS^{t}$ subsets contain significantly larger unit cells compared to the training set (see Appendix E for statistics of their sizes), yet our method still excels on these subsets as measured by the multiple MAE metrics, demonstrating its good scalability on the sizes of atomic systems it handles.
@@
In Table 2, we present the results of QHNet and QHNet+TraceGrad under the metrics of $MAE^{H}_{all}$, $MAE^{H}_{diag}$, $MAE^{H}_{non\_diag}$, $MAE^{\epsilon}$, and $Sim(\psi)$ for the $QS$ and $QD$ databases. The results of QHNet to be compared are taken from their original paper (Yu et al., 2023a), and for the unification of MAE units, we convert the units of MAE from $10^{-6}$ Hartree ($E_{h}$) in the original paper to meV ($27211.41E_{h}=$ 27211.4 meV). The results of QHNet+TraceGrad are the average from 10 independent repeated experiments. To ensure reproducibility, we use the same fixed random seeds as employed in QHNet for all random processes in the experiments on the $QS$ and $QD$ databases. As a result, for each Hamiltonian MAE result presented in Table 2, the standard deviation across repeated experiments is no greater than $0.009$ meV and is also negligible.
@@
The results presented in Table 2 demonstrate that the proposed TraceGrad method significantly enhances the accuracy of the baseline QHNet model across all metrics on the $QS$ and $QD$ databases. Specifically, TraceGrad reduces $MAE^{H}_{all}$, $MAE^{H}_{non\_diag}$, $MAE^{H}_{diag}$, $MAE^{\epsilon}$, and $Sim(\psi)$ of QHNet with relative reductions by a maximum of 40%, 39%, 40%, 76%, and 17%, respectively, on the corresponding databases. The significant accuracy improvements on the $QS$ database, partitioned using the ’ood’ split strategy (Yu et al., 2023a) without scale overlapping among the training, validation, and testing sets, once again demonstrate the method’s strong generalization capabilities across different scales of atomic systems. Meanwhile, performance on the $QD$ database under the ’mol’ strategy (Yu et al., 2023a), which partitions the training, validation, and testing sets with samples from completely different thermal motion trajectories, highlight our method’s robustness in generalizing to new thermal motion sequences. Furthermore, the substantial improvement in the prediction accuracy of $\epsilon$, i.e., occupied orbital energies crucial for determining electronic properties such as optical characteristics and conductivity in atomic systems, and $\psi$, i.e., the electronic wavefunctions essential for understanding electron distribution and interactions, underscores the potential values of our method for applications like material design, molecular pharmacology, and quantum computing.
We also evaluate the acceleration performance of our method for DFT computations and report the results in Appendix K . Experimental results show that while combining TraceGrad introduces only a slight increase from the inference time of QHNet, it delivers significant improvements in accelerating the convergence of DFT methods.
7 Summary of Appendices
Due to the page limit of the main manuscript, we have to provide some important contents in the Appendices, summarized as follows:
Appendix A provides the definitions of foundational concepts relevant to the problem addressed in this paper; Appendix B describes the application task in this paper, i.e., the electronic-structure Hamiltonian prediction task; Appendix C formalizes Hamiltonian prediction as an SO(3)-equivariant non-linear representation learning problem; Appendix D provides the proofs for all of the proposed theorems; Appendix E presents the detailed information of the experimental databases; Appendix F presents the implementation details of the experiments; Appendix G compares the block-level MAE statistics ($MAE^{H}_{block}$ and $MAE^{H}_{cha\_b}$) for DeepH-E3 and DeepH-E3+TraceGrad; Appendix H reports the results of ablation study for each individual mechanism of our framework and the quantitative comparison between the proposed gradient-based mechanism (Grad) with the existing gated activation mechanism (Gate). Experimental results indicate that each individual mechanism of our method can contribute individually to the performance. Moreover, their combination provides even better performance. Experimental results also demonstrate better accuracy performance of the proposed gradient-based mechanism compared to the gated mechanism; Appendix I provides a theoretical analysis of the computational complexity advantage of our method over traditional DFT calculations; Appendix J makes a joint discussion on GPU time costs and performance gains brought by our TraceGrad method. It underscores the superiority of the TraceGrad method in improving accuracy performance while maintaining time efficiency; Appendix K reports the acceleration performance of our method for traditional DFT algorithms; Appendix L corresponds to the synergy of our method with an approximately SO(3)- equivariant methodology (Yin et al., 2024).
@@ 8 Conclusion
We propose a theoretical and methodological framework to tackle the issue of reconciling powerful non-linear expressiveness with SO(3)-equivariance in deep learning paradigm for electronic-structure Hamiltonian prediction, through deeply investigating the mathematical connections between SO(3)-invariant and SO(3)-equivariant quantities, as well as their representations. We first constructs SO(3)-invariant quantities from SO(3)-equivariant regression targets, using them to train informative SO(3)-invariant non-linear representations. From these, SO(3)-equivariant features are derived with gradient operations, achieving non-linear expressiveness while maintaining strict SO(3)-equivariance and physical dimensional conistency to the target labels. Experimental results on eight benchmark databases covering multiple types of systems and challenging scenarios show substantial improvements on the state-of-the-art prediction accuracy of deep learning paradigm on electronic-structure Hamiltonian prediction. Our method also significantly promotes the acceleration performance for the convergence of traditional DFT methods.
@@
A Definition of Concepts
This section provides definitions of foundational concepts relevant to the problem addressed in this paper. For additional background, please refer to the book by Dresselhaus et al. (2007).
Definition .1.
Group. A set $G$, denoted as $G=\{\dots,g,\dots\}$, equipped with a binary operation $\cdot$, is called a group if it satisfies the following four axioms:
@@ A Definition of Concepts
    Closure: For any $f,g\in G$, the result of their operation $f\cdot g$ is also an element of $G$: $f\cdot g\in G$.
@@
    Associativity: For all $f,g,h\in G$, the operation satisfies $(f\cdot g)\cdot h=f\cdot(g\cdot h)$.
@@
    Identity Element: There exists a unique element $e\in G$ (called the identity) such that for all $f\in G$, $e\cdot f=f\cdot e=f$.
@@
    Inverse Element: For each $f\in G$, there exists a unique element $f^{-1}\in G$ (called the inverse) such that $f\cdot f^{-1}=f^{-1}\cdot f=e$.
Definition .2.
SO(3) Group. The special orthogonal group $\text{SO}(3)$ is the group of all $3\times 3$ orthogonal matrices with determinant 1. Formally, it is defined as:
$\text{SO}(3)=\{\mathbf{R}\in\mathbb{R}^{3\times 3}\mid\mathbf{R}^{T}\mathbf{R}=I,\,\det(\mathbf{R})=1\},$
where $\mathbf{R}^{T}$ denotes the transpose of $\mathbf{R}$, and $I$ is the $3\times 3$ identity matrix. The elements of $\text{SO}(3)$ represent rotations in three-dimensional Euclidean space.
Definition .3.
Group Representation. A representation of a group $G$ on a tensor space $T(V)$ is a homomorphism $\rho$ from $G$ to the general linear group $GL(T(V))$, the group of all invertible linear transformations on $T(V)$. Here, $T(V)$ represents the tensor space associated with the vector space $V$, encompassing all tensors that can be formed from elements of $V$. Formally, the homomorphism $\rho$ is defined as:
@@
Definition .4.
Irreducible Representation. A representation $\rho:G\to GL(V)$ of a group $G$ on a vector space $V$ is said to be irreducible if there is no proper subspace $W\subset V$ such that $\rho(g)W\subset W$ for all $g\in G$. In other words, the only invariant subspaces under the group action are the trivial subspaces $\{0\}$ and $V$ itself. If such a nontrivial invariant subspace exists, the representation is said to be reducible.
Definition .5.
SO(3) Group Representation. A representation of the special orthogonal group $\text{SO}(3)$ on a vector space $V$ is a homomorphism
$\rho:\text{SO}(3)\to GL(V),$
where $GL(V)$ denotes the group of all invertible linear transformations on $V$. The irreducible representation of $\text{SO}(3)$, each labeled by a degree $l$, often corresponds to the quantum number for angular momentum in quantum physics.
Definition .6.
Wigner-D Matrices. One of the irreducible representations of the rotation group $\text{SO}(3)$ is given by the Wigner-D matrices $\mathbf{D}^{l}(\mathbf{R})$, where $l$ is the degree of the representation, typically associated with angular momentum quantum number in quantum mechanics.
The Wigner-D matrices are defined as:
$D^{l}_{m^{\prime}m}(\mathbf{R})=\langle l,m^{\prime}|\mathbf{R}|l,m\rangle,$
where $|l,m\rangle$ are the eigenstates of the angular momentum operator $\hat{L}^{2}$ and its $z$-component $\hat{L}_{z}$, with $m$ and $m^{\prime}$ being the magnetic quantum numbers corresponding to the eigenvalue of $\hat{L}_{z}$. The Wigner-D matrices $\mathbf{D}^{l}(\mathbf{R})$ encode how the angular momentum states $|l,m\rangle$ transform under the action of a rotation.
Definition .7.